OpenStudy (vianne):

3 3/4 x (-1 1/2) = ?

OpenStudy (anonymous):

3 3/4 x (-1 1/2) = 15/4 × -3/2 = (15 × -3) / (4 × 2) = -45/8 = -45/8 = -5 5/8

OpenStudy (vianne):

What are the ��� for? I couldn't understand most of the equation because of the ��� in the way on my screen..